HELPPPP!!!
solve each equation below.
the worksheet is titled: double cross.
g. x/2+2x/3=5
t. 5x/6-(3-x)/8=(4x+3)/12
s. 1/3(a+5)=7/2
b. (3k-8)/14+5/7=(k+1)/2

key: / means divide.
Sorry i wish i could find out a way to write the divide sign as it is on the paper i have so it would be easier to understand.
I am trying to help my little brother with his math and i am stuck on these problems!! PLEASE PLEASE help... he needs it badly!! help us!!
THANK YOU!!!!!!!!!!!!

Answer by Theo(2967)

g. x/2+2x/3=5
x/2 + 2x/3 = 5
multiply both sides of the equation by 6 to get:
3x + 4x = 30
combine like terms to get:
7x = 30
divide both sides of the equation by 7 to get:
x = 30/7

substitute in original equation to get:

x/2 + 2x/3 = 5 becomes:

(30/7)/2 + (2*(30/7))/3 = 5

multiply both sides of the equation by 6 to get:

3*(30/7) + 2*(2*(30/7)) = 6*5

simplify to get:

90/7 + 120/7 = 30

combine like terms to get:

210/7 = 30

simplify to get:

30 = 30 confirming that x = 30/7 is good.

It's easiest to remove fractions by multiplying both sides of the equation by a number that will eliminate them.

Just be careful to multiply both sides of the equation by the same amount to preserve the equality.

One little slip messes the whole thing up.

----------------------------------------------

t. 5x/6-(3-x)/8=(4x+3)/12

I think this is:

(5x/6) - (3-x)/8 = (4x+3)/12

multiply both sides of the equation by 12 to get:
10x - (12*(3-x))/8 = 4x+3

multiply both sides of the equation by 8 to get:

80x - 12*(3-x) = 8*(4x+3)

simplify by removing parentheses to get:

80x - 36 + 12x = 32x + 24

combine like terms to get:

92x - 36 = 32x + 24

subtract 32x from both sides of the equation to get:

92x - 32x - 36 = 24

add 36 to both sides of the equation to get:

92x - 32x = 24 + 36

combine like terms to get:

60x = 60

divide both sides of the equation by 60 to get:

x = 1

substitute in the original equation to get:

(5x/6) - (3-x)/8 = (4x+3)/12 becomes:

(5/6) - (2/8) = (7/12)

multiply both sides of the equation by 12 to get:

10 - (24/8) = 7

simplify to get:

10 - 3 = 7

combine like terms to get:

7 = 7 confirming the value of x = 1 is good.

---------------------------------------

s. 1/3(a+5)=7/2

multiply both sides of the equation by 3 to get:

a + 5 = 21/2

multiply both sides of the equation by 2 to get:

2a + 10 = 21

subtract 10 from both sides of the equation to get:

2a = 11

divide both sides of the equation by 2 to get:

a = 11/2

substitute in the original equation to get:

1/3(a+5)=7/2 becomes:

(1/3)*((11/2) + 5) = 7/2

Multiply both sides of the equation by 6 to get:

6*(1/3)*((11/2)+5) = 6*7/2

simplify to get:

2*((11/2+5) = 3*7

simplify further to get:

11 + 10 = 21

combine like terms to get:

21 = 21 confirming the value of a = 11/2 is good.

---------------------------------

b. (3k-8)/14+5/7=(k+1)/2

I think this is:

((3k-8)/14) + (5/7) = (k+1)/2

assuming that it is, then the equation would be solved as follows:

multiply both sides of the equation by 14 to get:

(3k-8) + 10 = 7*(k+1)

simplify by removing parentheses to get:

3k-8+10 = 7k+7

combine like terms to get:

3k+2 = 7k+7

subtract 3k from both sides of the equation to get:

2 = 7k+7-3k

subtract 7 from both sides of the equation to get:

2-7 = 7k-3k

combine like terms to get:

-5 = 4k

this is the same as:

4k = -5

divide both sides of the equation by 4 to get:

k = -5/4

substitute in the original equation to get:

(3k-8)/14) + (5/7) = (k+1)/2 becomes:

(3*(-5/4) -8)/14 + (5/7) = ((-5/4)+1)/2

Simplify to get:

((-15/4)-8)/14 + (5/7) = (-1/4)/2

multiply both sides of the equation by 14 to get:

((-15/4) -8) + 10 = 7*(-1/4)

remove parentheses to get:

(-15/4) - 8 + 10 = (-7/4)

Multiply both sides of the equation by 4 to get:

-15 - 32 + 40 = -7

combine like terms to get:

-7 = -7 confirming that x = -5/4 is a good answer.

-----------------------------------------